OpenStudy (anonymous):

What value is added to both sides of the equation x2 - 8x = -10 in order to solve by completing the square? -8 16 -4 4

zepdrix (zepdrix):

\[\Large\rm x^2+\color{orangered}{b}x=stuff\]To complete the square, we need to find that special value that we're adding to each side. Take half of the b term and square it.\[\Large\rm x^2+\color{orangered}{(-8)}x=-10\]

OpenStudy (anonymous):

16

zepdrix (zepdrix):

\[\Large\rm \left(\frac{\color{orangered}{b}}{2}\right)^2\]\[\Large\rm \left(\frac{\color{orangered}{-8}}{2}\right)^2\]Ooo yay good job \c:/